Shift report 20000824 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 24 2000 - 00:02:06 EDT

  • Next message: http account: "Shift report 20000824 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 24 00:02:06 2000, Shift: 16:00-24:00
    Supervisor: FR, with: ZM and RD
    Log book#5 pages 3 - 4
    
    Shift Summary:
    ------------------------------------------------------------
     1. Beam was cogged at 20:55 with I_b=15.4e9 and I_y=12.1e9.
      Trigger 1 rate was around 120Hz at the beginning of the 
      store.      
     2. Data taken with:
       - MRS at 90deg., D5 on (350A, pol A).
       - FFS at 8deg., D1 on (625A, pol B), D2 on (492A, pol B).
       - T1 was off
     Run 4221: about 75 min., contains 220K trigger 1 events.
     Run 4222: started at 22:15, with T1 turned on, run still
     going.
     3. No new detector problems.
    
     The Forward Spectrometer has been moved from 5 to 8 degrees
     at the beginning of the shift.
     KO included in the "Begin run" of DAQ panel new entry
     fields for the spectrometer angles and magnet settings
     and polarities.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 24 2000 - 00:05:29 EDT